Zsazsa is twice as old as Mayet. The sum of their ages is 4 times Mayet's age less 46.How old are Zsazsa and Mayet?
Answer by Maths68(1474)

Let
Present age of Zsazsa = z years old
Present age of Mayet = m years old

Zsazsa is twice as old as Mayet
z=2m...........(1)
The sum of their ages is 4 times Mayet's age less 46
z+m=4m-46............(2)
Put the value of z from (1) to (2)
z+m=4m-46............(2)
2m+m=4m-46
3m=4m-46
3m-4m=-46
-m=-46
Multiply by -1 both sides
m=46
Put the value of m in (1)
z=2m
z=2(46)
z=92


Present age of Zsazsa = z = 92 years old
Present age of Mayet = m = 46 years old